OpenStudy (anonymous):

find the product. Simplify your answer. (8x)/(-1) · (1)/(4x^2)

OpenStudy (anonymous):

-32X^3

OpenStudy (anonymous):

nm thats division

OpenStudy (anonymous):

rational expressions really aren't my thing

OpenStudy (anonymous):

\[\frac{ a }{ b } \times \frac{ c }{ d } = \frac{ ac }{ bd }\]

OpenStudy (anonymous):

would it be \[2/-x\]

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

thank you